OpenStudy (anonymous):

Simplify the following expression \[\sqrt{-90}\sqrt{-15}\]

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

i'd help you if i could but i don't know anything about complex numbers, and you'd have to use those right? :P

OpenStudy (anonymous):

DO YOU KNOW HOW TO DO SCIENTIFIC NOTATION NUMBER

OpenStudy (anonymous):

yeah. the answer to your question though is -15(6)^0.5

OpenStudy (anonymous):

DO YOU UNDERSTAND HOW TO DO Use scientific notation to divide the following two numbers. Express the answer using scientific notation; retain at least three decimal places. (1.83 · 106) / (1.2 · 104)

OpenStudy (anonymous):

any help anyone

OpenStudy (unklerhaukus):

complex numbers are not required to solve this problem \[\sqrt{-90}\sqrt{-15}\]\[=\sqrt{(-90 \times -15)} \]\[=\sqrt{-1\times 3^2\times2\times5\times -1\times 3\times 5}\]\[=\sqrt{-1^2 \times 2 \times 3^3\times 5^2}\]\[=\sqrt{2\times3^3\times5^2}\]\[=5\times 3\sqrt{2\times3}\]\[=15\sqrt6\]\[\approx36.74=3.674\times10^1\]

OpenStudy (unklerhaukus):

\[\frac{(1.83 \times 10^6) }{ (1.2 \times 10^4)}\]\[=\frac{1.83}{1.2}\times10^2\]\[=1.525\times10^2\]

OpenStudy (anonymous):

complex number are required...